i need help with the following please:
Determine which two equations represent perpendicular lines.
(a) y = 5/8 x – 3
(b) y = 3x – 5/8
(c) y = -1/3 x + 5/8
(d) y = 1/3 x – 5/8



Answer by malakumar_kos@yahoo.com(315)


Determine which two equations represent perpendicular lines.
(a) y = 5/8 x – 3
(b) y = 3x – 5/8
(c) y = -1/3 x + 5/8
(d) y = 1/3 x – 5/8

The condition for the two lines to be perpendicular is
product of their slopes = -1
From th above given eq'ns (b) and (c) their slopes are 3 and -1/3
hence their product = 3*-1/3 = -1 which satisfies the condition
(b) and (c) are perpendicular to each other